# Low-Level and NUMA-Aware Optimization for High-Performance Quantum Simulation

### v1

@@ III-B Quantum Fourier Transform
To evaluate our approach on a more complex multi-qubit algorithm, we examined the QFT circuit. Figure 8(a) compares the execution times of the baseline QFT across ten qubit sizes using four different implementations: a divide-and-conquer recursive variant inspired by libraries like Yao/Qiskit, an iterative method, QuEST Built-in, and an optimized cache-blocked version with two blocking factors. The recursive implementation mirrors the mathematical definition of the QFT by recursively applying a Hadamard gate followed by controlled phase shifts. The iterative method, on the other hand, rearranges the qubit operations into nested loops instead of relying on recursion. The iterative layout can be more cache-friendly in certain scenarios due to its straightforward memory access pattern, representing differences of less than 0.5–1.5% relative to the recursive approach across all qubit sizes. Although the built-in method is robust, it yields consistently higher execution times. At 23 qubits, the QuEST-Builtin runs 1.86× slower than the recursive/iterative variants, and at 32 qubits it’s roughly 1.70× slower. This indicates that the hand-optimized approaches can provide a significant performance gain – roughly a 70–80% reduction in runtime compared to the default library implementation. The QFT can be cache-blocked without extra gates by shifting its ending SWAP gates to the left, ensuring that all Hadamard gates operate on local qubits. Only the distributed SWAP gates (which are flipped vertically) incur communication overhead [15]. In an effort to improve data locality, we implemented a cache-blocking technique with a blocking factor of 16, which is 3% slower than the iterative approach at 23 qubits. At the upper end (32 qubits), the blocked variant is about 4.8% slower than the iterative version. These results suggest that for the chosen block size, the overhead associated with managing cache blocks slightly outweighs the potential gains in memory locality, particularly as the data size grows. When we increased the block size to 64, performance remained virtually identical to the iterative baseline – evidence that the iterative method already exploits strong memory locality. The larger block size simply further amortizes data‐movement costs and maximizes cache utilization without adding significant overhead.
@@
To evaluate the effectiveness of our optimizations, we compare two implementations of QFT – recursive and QuEST Built-in – as summarized in Table II. Overall, the recursive variant consistently achieves higher speedups (1.8x) than the Built-in (1.7x), indicating that its code structure may benefit more from compiler‐driven optimizations like inlining or function‐call elimination, giving it a minor advantage. Nevertheless, both implementations show robust improvements over their respective baselines due to low‐level enhancements. Hence, while both implementations benefit from our optimizations, the recursive QFT’s design appears to leverage these techniques slightly more effectively. Nonetheless, both QFT implementations are less amenable to extreme speedups compared to simple gate operations. The QFT’s series of dependent rotations and swaps impose a more sequential data access pattern, which limits the gains achievable from SIMD and caching techniques.
| Qubit | Recursive QFT | Built-in QFT |
@@ III-C Random Quantum Circuits
As a final benchmark, we implemented an RQC generator to stress-test the simulator. This setup allows us to quantify the impact of our low-level optimizations on both the dynamic generation of random single‑qubit gate sequences and the highly entangled execution of CNOTs connecting qubit →. Here, each circuit comprises a fixed number of layers (), and we sweep the register size from 23 to 32 qubits. We measured the performance of executing these circuits end-to-end, including state initialization, random gate generation, and state destruction, using the Google Benchmark framework for rigorous timing. Figure 9 presents the results for the baseline and optimized RQC executions. Despite the randomness of the circuits, their structure is fairly regular across qubits and layers, and our optimized simulator achieves a consistent speedup of 4× over the baseline. The RQC’s uniform layering and parallel gate structure are extremely well-suited to our low-level optimizations, which greatly reduce cache misses and overlap memory accesses with computation, thereby yielding a substantial boost in overall simulation speed. By contrast, as discussed above, the QFT’s inherently sequential structure and greater arithmetic complexity, which is primarily due to the interdependent controlled phase rotations introduces bottlenecks that constrain the effectiveness of SIMD strategies. Loop unrolling and prefetching do not capitalize on these operations as efficiently, and the sequential dependencies introduce communication overhead that further hinders performance improvements-limiting its speedup to 1.8×. This comparison highlights that the effectiveness of our optimizations can depend on the circuit’s structure – highly parallel circuits like RQCs reap the most benefit, whereas circuits with serial dependencies see more moderate gains.
@@ IV Discussion
Scaling. Our experiments only include configurations with 2 NUMA nodes. While this is not necessarily the most NUMA regions that can be observed in modern machines, we expect our evaluation to cover an overwhelming amount of them nonetheless, for two main reasons. Firstly, when more compute power is needed, modern systems often reach for CPUs with more cores per socket, or accelerators such as GPUs or TPUs, rather than large socket counts per node. Machines that need a large number of CPUs or GPUs usually fall directly under the distributed computing paradigm, where shared memory is no longer applicable [21]. Secondly, in the presence of several NUMA regions on 2 sockets, the penalties paid for going across sockets are much larger than when going across memory controllers on the same socket. We observe for example on ARCHER2 [22], a machine running dual-socket AMD® EPYC 7742 processors nodes with 8 NUMA regions (4 per processor), that going across sockets in the same node costs a 30% decrease in memory bandwidth – similar to the performance of the system we used for benchmarking – while going across NUMA regions on the same socket costs 2% to 5%. This loss is so much smaller than going across sockets, that functionally this 8-region machine behaves like a 2-region configuration in practice.
@@ V Conclusion
Our work bridges the gap between high-level algorithmic efficiency and low-level, architecture-specific performance enhancements. We have significantly enhanced the robust foundation of the QuEST simulator (v3.7) by developing a NUMA-aware, vectorized execution model optimized explicitly for modern multi-core processors. By restructuring the data layout to an AoS and applying targeted optimizations – including NUMA-aware memory allocation, explicit thread pinning, AVX-512 vectorization with intrinsic instructions, aggressive loop unrolling, and explicit memory prefetching – we achieved substantial speedups: 5.5–6.5× for single-qubit gates, 4.5× for two-qubit gates, 4× for RQCs, and approximately 1.8× for QFT circuits over QuEST. Our optimizations compare favorably against related works such as HpQC and PAS. HpQC similarly utilizes AoS memory layout and AVX2 intrinsics, reporting QFT speedups of 2.20× and RQC speedups around 1.74×. PAS incorporates distinct optimizations such as hybrid vectorization (combining intrinsic and compiler auto-vectorization), fast bitwise arithmetic, quantum tracking to avoid redundant computations, and memory access filtering, achieving higher performance gains – up to 8.69× on QFT and 2.62× on RQC. PAS also explicitly emphasizes large 4K-byte aligned memory allocations compared to our 64-byte alignment strategy, reflecting a more aggressive optimization for large memory pages. Despite these notable techniques in PAS and HpQC, our simulator uniquely integrates NUMA-aware memory placement and explicit core affinity, significantly reducing inter-node latency and better exploiting memory bandwidth on NUMA systems – features notably absent from PAS and HpQC. Moreover, unlike PAS and HpQC – which currently do not provide open access to their implementations – our enhancements are fully open-sourced, promoting transparency, reproducibility, and community-driven improvements. Our extension is hosted in our group’s repository [23], available upon reasonable request, and will soon be integrated into mainstream QuEST following completion of integration testing. This will enable a comprehensive benchmarking of our optimizations directly against the latest QuEST v4 release [24], which introduced new algorithmic methods [2] and architectural enhancements, but without incorporating our detailed NUMA-aware and SIMD-based tuning proposed here. Such benchmarking will quantitatively assess the specific performance impacts attributable to our low-level optimizations. Ultimately, our framework provides a modular, extensible foundation for accelerating classical quantum simulations.
@@
Future Work. Going forward, we aim to maximize single-node performance and scalability through automation of vectorized kernel generation and compile-time-driven prefetching strategies, while extending support to larger shared-memory systems and heterogeneous accelerator-based platforms. In the near term, we plan to adapt NUMA-aware optimization concepts to GPUs, leveraging their shared emphasis on data locality. Although NUMA originated in CPU architectures, its core principles – memory locality and task affinity – are increasingly applicable to GPUs and can inform performance-aware execution across heterogeneous platforms [25]. Next, we will explore how emerging CXL memory expanders – i.e., CPU-less NUMA nodes impact the performance and applicability of the techniques introduced in this work. While CXL clearly extends the memory capacity of a single node, its effect on simulation performance and the strategies required to optimize it remain open questions. Finally, we will explore low-level and locality-sensitive challenges in distributed multi-node, multi-GPU environments that leverage MPI and RDMA for communication.

### v2

@@ III-B Quantum Fourier Transform
To evaluate our approach on a more complex multi-qubit algorithm, we examined the QFT circuit. Figure 8 compares the execution times of the baseline QFT across ten qubit sizes using four different implementations: a divide-and-conquer recursive variant inspired by libraries like Yao/Qiskit, an iterative method, QuEST Built-in, and an optimized cache-blocked version with two blocking factors. The recursive implementation mirrors the mathematical definition of the QFT by recursively applying a Hadamard gate followed by controlled phase shifts. The iterative method, on the other hand, rearranges the qubit operations into nested loops instead of relying on recursion. The iterative layout can be more cache-friendly in certain scenarios due to its straightforward memory access pattern, representing differences of less than 0.5–1.5% relative to the recursive approach across all qubit sizes. Although the built-in method is robust, it yields consistently higher execution times. At 23 qubits, the QuEST-Builtin runs 1.86 slower than the recursive/iterative variants, and at 32 qubits it is roughly 1.70 slower. This indicates that the hand-optimized approaches can provide a significant performance gain–roughly a 70–80% reduction in runtime compared to the default library implementation. The QFT can be cache-blocked without extra gates by shifting its ending SWAP gates to the left, ensuring that all Hadamard gates operate on local qubits. Only the distributed SWAP gates (which are flipped vertically) incur communication overhead [15]. In an effort to improve data locality, we implemented a cache-blocking technique with a blocking factor of 16, which is 3% slower than the iterative approach at 23 qubits. At the upper end (32 qubits), the blocked variant is about 4.8% slower than the iterative version. These results suggest that for the chosen block size, the overhead associated with managing cache blocks slightly outweighs the potential gains in memory locality, particularly as the data size grows. When we increased the block size to 64, performance remained virtually identical to the iterative baseline–evidence that the iterative method already exploits strong memory locality. The larger block size simply further amortizes data‐movement costs and maximizes cache utilization without adding significant overhead.
@@
To evaluate the effectiveness of our optimizations, we compare two implementations of QFT–recursive and QuEST Built-in–as summarized in Table II. Overall, the recursive variant consistently achieves higher speedups (1.8) than the Built-in (1.7), indicating that its code structure may benefit more from compiler‐driven optimizations like inlining or function‐call elimination, giving it a minor advantage. Nevertheless, both implementations show robust improvements over their respective baselines due to low‐level enhancements. Hence, while both implementations benefit from our optimizations, the recursive QFT’s design appears to leverage these techniques slightly more effectively. Nonetheless, both QFT implementations are less amenable to extreme speedups compared to simple gate operations. The QFT’s series of dependent rotations and swaps impose a more sequential data access pattern, which limits the gains achievable from SIMD and caching techniques.
| Qubit | Recursive | Built-in |
@@ III-C Random Quantum Circuits
Next, we implemented an RQC generator to stress-test the simulator. This setup allows us to quantify the impact of our low-level optimizations on both the dynamic generation of random single‑qubit gate sequences and the highly entangled execution of CNOTs connecting qubit →. Here, each circuit comprises a fixed number of layers (), and we sweep the register size from 23 to 32 qubits. We measured the performance of executing these circuits end-to-end, including state initialization, random gate generation, and state destruction, using the Google Benchmark framework for rigorous timing. Figure 9 presents the results for the baseline and optimized RQC executions. Despite the randomness of the circuits, their structure is fairly regular across qubits and layers, and our optimized simulator achieves a consistent speedup of 4 over the baseline. The RQC’s uniform layering and parallel gate structure are extremely well-suited to our low-level optimizations, which greatly reduce cache misses and overlap memory accesses with computation, thereby yielding a substantial boost in overall simulation speed. By contrast, as discussed above, the QFT’s inherently sequential structure and greater arithmetic complexity, which is primarily due to the interdependent controlled phase rotations introduces bottlenecks that constrain the effectiveness of SIMD strategies. Loop unrolling and prefetching do not capitalize on these operations as efficiently, and the sequential dependencies introduce communication overhead that further hinders performance improvements-limiting its speedup to 1.8. This comparison highlights that the effectiveness of our optimizations can depend on the circuit’s structure–highly parallel circuits like RQCs reap the most benefit, whereas circuits with serial dependencies see more moderate gains.
QFT is primarily compute-bound and particularly sensitive to vectorization, while RQC is memory-bound and strongly exposes data layout and NUMA effects. This complementary pairing has been used in prior simulation studies to isolate the effects of different architectural and kernel-level optimizations [20, 12, 4]. However, these two workloads alone do not fully capture the behavior of structured algorithmic circuits. To broaden the evaluation scope, we additionally consider Grover and Shor-like periodicity estimation circuits, which introduce distinct entanglement and control patterns and offer a more representative view of end-to-end performance.
III-D Grover and Shor-like Workloads
The execution time comparison presented in Figure 10 and the corresponding speedup in Table III clearly illustrate the impact of the optimized kernel on two representative circuit-level workloads, Grover and Shor. For both circuits, execution time increases exponentially with the number of qubits, as expected from the underlying state vector scaling. However, the optimized implementation achieves a substantial reduction in runtime across the entire range of qubit counts, reflecting the combined effect of NUMA-aware scheduling, AVX-512 vectorization, and cache-aware prefetching. The gains are particularly evident at higher qubit numbers, where memory bandwidth becomes the dominant bottleneck.
For Grover, which has relatively shallow depth and lower gate density, the optimized kernel delivers a stable speedup of approximately – from 23 to 32 qubits. This indicates that the vectorized Hadamard, prefetching strategy, and NUMA-aware scheduling effectively sustain high memory throughput across all problem sizes. For the Shor-like circuit, which is substantially more gate-dense and involves controlled operations and an inverse QFT, the achieved speedups are lower but still significant, consistently around across the entire range. This reflects the heavier synchronization and memory traffic inherent in Shor, while still demonstrating a clear performance gain from the optimized kernels. These results confirm that the applied optimizations scale robustly for both low-depth and highly entangling algorithmic workloads.
| Qubit | Grover Speed-up | Shor Speed-up |
| 23 | 4.58 | 2.57 |
| 24 | 4.50 | 2.62 |
| 25 | 4.46 | 2.51 |
| 26 | 4.46 | 2.51 |
| 27 | 4.49 | 2.54 |
| 28 | 4.50 | 2.52 |
| 29 | 4.49 | 2.55 |
| 30 | 4.48 | 2.54 |
| 31 | 4.50 | 2.50 |
| 32 | 4.34 | 2.56 |
@@ IV Discussion
Scaling. Our experiments only include configurations with 2 NUMA nodes. While this does not represent the maximum number of NUMA regions found in modern machines, we expect our evaluation to cover most practical configurations of them nonetheless, for two main reasons. Firstly, when more compute power is needed, modern systems often reach for CPUs with more cores per socket, or accelerators such as GPUs or TPUs, rather than large socket counts per node. Machines that need a large number of CPUs or GPUs usually fall directly under the distributed computing paradigm, where shared memory is no longer applicable [23]. Secondly, in the presence of several NUMA regions on 2 sockets, the penalties paid for going across sockets are much larger than when going across memory controllers on the same socket. We observe for example on ARCHER2 [24], a machine running dual-socket AMD® EPYC 7742 processors nodes with 8 NUMA regions (4 per processor), that going across sockets in the same node costs a 30% decrease in memory bandwidth–similar to the performance of the system we used for benchmarking–while going across NUMA regions on the same socket costs 2% to 5%. This loss is so much smaller than going across sockets, that functionally this 8-region machine behaves like a 2-region configuration in practice.
@@ V Conclusion
Our work bridges the gap between high-level algorithmic efficiency and low-level, architecture-specific performance enhancements in classical quantum simulation. We significantly extend the robust foundation of the QuEST simulator (v3.7.0) by developing a NUMA-aware, vectorized execution model explicitly optimized for modern multi-core processors. By restructuring the state vector to an AoS layout and applying a suite of targeted optimizations–including NUMA-aware memory allocation, explicit thread pinning, AVX-512 vectorization through intrinsic instructions, aggressive loop unrolling, and explicit memory prefetching–we achieved substantial and consistent speedups across a range of workloads. For low-level primitives, we obtained speedups of – for single-qubit gates and for two-qubit gates across 23–35 qubits. For circuit-level benchmarks, the optimized kernels delivered a speedup on RQC and approximately on QFT circuits. Extending beyond these benchmarks, we further evaluated two representative algorithmic workloads: Grover’s search and a Shor-like periodicity estimation circuit. These structured circuits, involving different entanglement and control patterns, exhibited stable performance gains of approximately – for Grover and for Shor across 23–32 qubits, confirming the generality and robustness of the optimizations. Our results compare favorably against related efforts such as HpQC and PAS. HpQC similarly employs AoS and AVX2 vectorization, reporting speedups of on QFT and on RQC, while PAS incorporates hybrid vectorization and fast bitwise arithmetic, reaching on QFT and on RQC. Unlike these approaches, our simulator integrates explicit NUMA-aware memory placement and thread affinity, substantially reducing inter-node latency and improving bandwidth utilization–features absent in both HpQC and PAS. Furthermore, while PAS and HpQC are not openly available, our implementation is fully open-sourced, promoting transparency, reproducibility, and community-driven improvements. The codebase is hosted in our group’s repository [25] and will be integrated into mainstream QuEST following completion of compatibility and performance testing. This integration will enable a direct, quantitative comparison of our optimizations against the latest QuEST v4.0 release [26], which introduced new algorithmic features but not the low-level NUMA-aware and SIMD-based tuning presented here. In summary, this work provides a modular and extensible optimization framework that yields consistent performance gains across diverse workloads–from primitive gates and random circuits to structured algorithmic circuits such as Grover and Shor. These results highlight the importance of co-optimizing memory locality, vectorization, and thread placement in accelerating classical quantum simulations on modern multi-socket architectures.
@@
Future Work. Going forward, we aim to maximize single-node performance and scalability through automation of vectorized kernel generation and compile-time-driven prefetching strategies, while extending support to larger shared-memory systems and heterogeneous accelerator-based platforms. In the near term, we plan to adapt NUMA-aware optimization concepts to GPUs, leveraging their shared emphasis on data locality. Although NUMA originated in CPU architectures, its core principles–memory locality and task affinity–are increasingly applicable to GPUs and can inform performance-aware execution across heterogeneous platforms [27]. Next, we will explore how emerging CXL memory expanders–i.e., CPU-less NUMA nodes–impact the performance and applicability of the techniques introduced in this work. While CXL clearly extends the memory capacity of a single node, its effect on simulation performance and the strategies required to optimize it remain open questions. Finally, we will explore low-level and locality-sensitive challenges in distributed multi-node, multi-GPU environments that leverage MPI and RDMA for communication.